OpenStudy (anonymous):

What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Let's think about this. That sector is what part of the entire circle?

OpenStudy (anonymous):

Alright? so

OpenStudy (anonymous):

The formula for the ara of a sector is n/360 x (pi)(radius)^2, where n = number of degrees in the central angle of the arc. In your case, n = 360; so the area of the sector will be 90/360 x (pi)(1)^2 = 1/4 x pi or pi/4.